OpenStudy (anonymous):

find two number with a sum of 5,256 and a difference of 2,142?

OpenStudy (anonymous):

x + y = 5256 x - y = 2142 Then subtract the equations from eachother to eliminate the X value (x+y) - (x - y) = 5256 - 2142 = 3114. Now, from there, your equation is 2y = 3114, making y=1557 Now, put that in your original equation (either one will work). 5256-1557= 3699. x= 3699. Test: 3699+1557= 5256 3699-1557= 2142

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem. Glad to help